# Grounding Language with Vision: A Conditional Mutual Information Calibrated Decoding Strategy for Reducing Hallucinations in LVLMs

### v1

@@ 1 Introduction
The unprecedented breakthroughs in large vision-language models (LVLMs) [1, 2, 3, 4] have expanded their applicability across various vision-language (V+L) tasks such as autonomous driving [5, 6]. Benefiting from advanced designs of model architectures and training algorithms, LVLMs trained on high-quality image-text pairs have exhibited outstanding capabilities in cross-modal alignment and complex V+L understanding. Despite the remarkable success, the issue of hallucination continues to pose challenges to LVLMs. Concretely, LVLMs may generate semantically coherent yet factually incorrect contents that are entirely inconsistent with the input image [7, 8, 9]. E.g., describe non-existent objects or misinterpret the attributes and relationships of visual entities within the image. This raises serious concerns regarding the deployment of LVLMs in real-world applications, particularly in high-risk scenarios such as medical diagnosis [10] and financial systems [11].
@@
To address this issue, we introduce conditional pointwise mutual information (C-PMI) to quantify the mutual correlation between the visual inputs and the generated texts during generation. Correspondingly, we reformulate the hallucination mitigation objective as a vision-language mutual information maximization problem, which is further decomposed into two complementary sub-tasks that capture the respective contributions of visual and textual tokens. Based on the analysis, we derive a bi-level optimization formulation and design an effective solution that adaptively calibrates each decoding step during the generation process. To optimize the inner sub-problem, we calibrate the token distribution using the derived formula to prioritize tokens that exhibit strong relevance to the visual input. For the outer sub-problem, we propose an efficient visual token purifier parameterized as a learnable network, to dynamically refine image tokens that are most pertinent to the current textual context. By filtering out redundant image tokens that impair mutual information with the generated content, the proposed strategy directs the model to focus more on the key visual tokens most relevant to the ongoing response (see Fig. 1 (b)), further enhancing the dependence of the generated text on the input image. To summarize, our main contributions are threefold:
@@ 2 Related Work
Large Vision-Language Models. Built upon advanced pre-trained LLMs [21, 22], LVLMs successfully bridge the gap between visual perception and linguistic reasoning [23, 24, 25, 4, 3, 26, 27], achieving impressive performance in generating diverse responses and tackling complex visual understanding tasks. To incorporate visual information into the LLM backbone, LVLMs like LLaVA [2, 4] and Shikra [3] employ linear projection layers trained by instruction fine-tuning to directly map visual features into the LLM embedding space. Meanwhile, the BLIP series [25, 28] introduces Q-former to integrate visual tokens dynamically through gated cross-attention layers, thereby reducing redundancy in image token representations. Benefiting from better training data, improved algorithms, and increasingly powerful LLM backbones, recent LVLMs such as LLaVA-Next [29] have demonstrated stronger multimodal understanding capabilities. Despite the progress, LVLMs still suffer from serious hallucination problems, where the generated responses are plausible yet unfaithful or factually incorrect. Our work aims to mitigate this issue and enhance the reliability of LVLMs.
@@ 3.1 Preliminary
Figure 2: Overview of the proposed multimodal decoding strategy at timestep $t$. To maximize the mutual relevance, CMI-VLD incorporates current image and text tokens as input to predict an image mask $\mathcal{M}_{v}$ for visual token refinement. Then, a token distribution based on Eq. (5) is correspondingly constructed for the next token sampling. For the training of the visual purifier, please refer to Sec. 3.3.
@@ 3.2 The Proposed CMI-VLD
(1) Calibrated Distribution Sampling for Text Modality. To optimize the text sequence $y$, Eq. (4) encourages us to construct an improved distribution $p_{c}$ to prioritize text tokens that maximize the difference between probabilities predicted with and without the visual input. However, directly applying this formulation yield unsatisfactory results since it can excessively penalize reasonable tokens in certain contexts. Inspired by [16, 8], we introduce a hyperparameter $\lambda$ to provide a more fine-grained control over the strength of the subtraction, which can be formally expressed as:
@@
(2) Visual Token Refinement for Visual Modality. Inspired by recent findings [34, 35] that many image tokens in LVLMs are redundant, we propose a visual token purification mechanism that enhances C-PMI by evicting tokens considered non-informative with respect to the given text. In this way, the LVLM can focus more on the most critical visual tokens for improved generation. Moreover, we also incorporate the model’s attention scores over the visual input to identify tokens that exert a stronger influence on the LVLM’s decisions. Specifically, given the query vectors $Q_{i}\in\mathbb{R}^{H\times n\times d_{k}}$ and key vectors $K_{i}\in\mathbb{R}^{H\times n\times d_{k}}$ at the $i$-th LVLM layer, where $H$ is the head number, $n$ is the current token number, and $d_{k}$ is the latent dimension, the total attention scores of an image $v$ is calculated as:
@@
where $\lambda$ is the aforementioned correction factor for distribution calibration and $\beta$ is a non-negative hyperparameter balancing the influence of C-PMI and attention scores on visual token selection.
At decoding step $t$, we first optimize the upper subproblem by sampling $y_{t}$ from the distribution adjusted based on Eq. (5). To solve the lower subproblem, we then adaptively retain a proportion $\gamma$ of image tokens as the purified input to promote its relevance to the current textual context. Motivated by findings in [35] that token sparsification at the second layer of LVLMs yields optimal performance, we utilize attention scores from this layer ($i=2$) for visual purification and start the refinement accordingly. By alternately solving the two subproblems at each decoding step, the optimizer simultaneously samples text tokens that tightly align with current visual content and purifies informative visual tokens with strong relevance to the ongoing textual context, hence effectively amplifying C-PMI and reducing hallucination-related elements in the final response.
@@ 3.3 Visual Token Purifier for Visual Refinement
To address the outer subproblem in Eq. (7), an intuitive solution is to manually select image tokens that maximize the defined score for every decoding step. However, this requires repeatedly calculating token-wise scores and would incur substantial computational burdens compared to existing decoding-based approaches [16, 17]. To overcome this challenge, we propose a lightweight visual token purifier $\mathcal{P}(\cdot)$, which consists of only a few transformer blocks and MLP layers (see Appendix C for details) [34], to automatically filter visual tokens that benefit C-PMI maximization.
@@ 4 Experiments
We conduct experiments on different LVLMs across various benchmarks to show the effectiveness of our method. For more experimental results and visualizations, please refer to Appendix D and F.
@@ 4.1 Experimental Setup
Models and Baselines. We align with [8] and choose four representative LVLMs for evaluation, including InstructBLIP [28], Shikra [3], LLaVA-1.5 on the 7B scale [4], and LLaVA-NeXT [29] on the 8B scale, which include two major categories of LVLMs that employ Q-former and linear layers as cross-modal connector respectively. All of these models adopt a pretrained visual encoder like CLIP [36], alongside a powerful LLM backbone such as LLaMA [21]. We conduct a comprehensive evaluation of the proposed CMI-VLD method on a range of state-of-the-art (SOTA) baselines, including Sampling (Top-p=1), Greedy, VTI [37], and LVLM decoding-based strategies, including VCD [16], ICD [17], HALC [7], OPERA [18], SID [8], and VASparse [9]. Following SID, we implement the proposed method under both sampling and greedy decoding settings. It is worth noting that HALC, OPERA, and VASparse adopt the more flexible and stronger beam search strategy, which may raise concerns regarding fairness. Nevertheless, extensive experiments demonstrate that our CMI-VLD still consistently outperforms these methods.
Implementation Details. We set $\alpha$ = $1\times 10^{2}$ and $\beta$ = $5\times 10^{2}$ in Eq. (9) during the training of the purifier for all LVLMs. For the layer to initiate the refinement process, we adopt Layer $i=2$ for LLaVA-1.5, Shikra, and LLaVA-NeXT and $i=4$ for InstructBLIP. Note that our method is compatible with the feature steering mechanism [37], which is then incorporated in our implementation. In Sec. 4.3, we conduct a detailed exploration of the strength of the subtraction $\lambda$ ranging from 00 to $0.9$. For baselines, we adopted the settings specified in the original papers or publicly available code of the baselines. For all the experiments, we adopt the widely used settings [18, 8] that set the max new tokens as 512 for evaluation. More implementation details are provided in Appendix B.
@@ 4.2 Performance Evaluation
POPE Evaluation. The POPE metric also focuses on object hallucinations by querying LVLMs to answer a yes/no question about whether a specific <object> is in the image. We report the results of the accuracy and F1 score in Table 2. The quantitative results reveal that our method generally performs best across the three split datasets. Notably, in the POPE evaluation, where responses are typically short and follow fixed patterns such as "Yes, there is a <object> in the image.", the evaluation primarily hinges on the first one or few tokens (i.e., Yes or No)  [18]. Consequently, our CMI-VLD may not fully exhibit its effectiveness in this constrained setup, as it is designed to dynamically adjust decoding over the entire generation rather than concentrating solely on the initial tokens. Nevertheless, our method still achieves notable improvements over competitive baselines. Due to the page limit, results on more LVLMs are provided in Appendix D.
MME and MMBench Evaluations. Apart from the above benchmarks tailored for hallucination evaluation, we additionally test on two popular LVLM benchmarks, i.e., MME [6] and MMBench [43], to systematically analyze their various capability dimensions. Following OPERA [18] and SID [8], we present results of the LLaVA-1.5 7B model as a representative in Table 3. As observed, CMI-VLD not only reduces the hallucinated contents but also enhances diverse capabilities of MLLMs, bringing remarkable improvements over the default decoding methods. These results underscore CMI-VLD as a reliable and practical strategy for hallucination mitigation.
Inference Time Analysis. Since our method introduces an additional visual token purifier for effective visual refinement, it is crucial to assess its influence on the overall computational efficiency. Following [19], we calculate the generation time per response to assess computing burdens. Figure 4 reveals that the proposed CMI-VLD achieves satisfactory decoding efficiency, introducing negligible computational overhead. This is primarily attributed to the lightweight architecture of the visual purifier and the removal of redundant visual tokens that would incur significant computational overhead, demonstrating that CMI-VLD effectively balances performance and efficiency.
4.3 Ablation Study
Next, we provide ablation analysis regarding several critical hyperparameters. More ablation analysis about the retaining ratio and the effectiveness of the proposed techniques is presented in Appendix D.
[…]

### v3

@@ 1 Introduction
The unprecedented breakthroughs in large vision-language models (LVLMs) [1, 2, 3, 4, 5] have expanded their applicability across various vision-language (V+L) tasks such as autonomous driving [6, 7]. Benefiting from advanced designs of model architectures and training algorithms, LVLMs trained on high-quality image-text pairs have exhibited outstanding capabilities in cross-modal alignment and complex V+L understanding. Despite the remarkable success, the issue of hallucination continues to pose challenges to LVLMs. Concretely, LVLMs may generate semantically coherent yet factually incorrect contents that are entirely inconsistent with the input image [8, 9, 10]. E.g., describe non-existent objects or misinterpret the attributes and relationships of visual entities within the image. This raises serious concerns regarding the deployment of LVLMs in real-world applications, particularly in high-risk scenarios such as medical diagnosis [11] and financial systems [12].
@@
To mitigate this issue, we introduce conditional pointwise mutual information (C-PMI) to quantify the mutual correlation between the visual inputs and the generated texts during generation. Correspondingly, we reformulate the hallucination mitigation objective as a vision-language mutual information maximization problem, which is further decomposed into two complementary sub-tasks that capture the respective contributions of visual and textual tokens. Based on the analysis, we derive a bi-level optimization formulation and design an effective solution that adaptively calibrates each decoding step during the generation process. To optimize the inner sub-problem, we calibrate the token distribution using the derived formula to prioritize tokens that exhibit strong relevance to the visual input. For the outer sub-problem, we propose an efficient visual token purifier parameterized as a learnable network, to dynamically refine image tokens that are most pertinent to the current textual context. By filtering out redundant image tokens that impair mutual information with the generated content, the proposed strategy directs the model to focus more on the key visual tokens most relevant to the ongoing response (see Fig. 1 (b)), further enhancing the dependence of the generated text on the input image. To summarize, our main contributions are threefold:
@@ 2 Related Work
Large Vision-Language Models. Built upon advanced pre-trained LLMs [22, 23, 24, 25], LVLMs successfully bridge the gap between visual perception and linguistic reasoning [26, 27, 3, 28, 29], achieving impressive performance in generating diverse responses and tackling complex visual understanding tasks. To incorporate visual information into the LLM backbone, LVLMs like LLaVA [2, 4] and Shikra [3] employ linear projection layers trained by instruction fine-tuning to directly map visual features into the LLM embedding space. Meanwhile, the BLIP series [30, 31] introduces Q-former to integrate visual tokens dynamically through gated cross-attention layers, thereby reducing redundancy in image token representations. Benefiting from better training data, improved algorithms, and increasingly powerful LLM backbones, recent LVLMs such as LLaVA-Next [32] have demonstrated stronger multimodal understanding capabilities. Despite the progress, LVLMs still suffer from serious hallucination problems, where the generated responses are plausible yet unfaithful or factually incorrect. Our work aims to mitigate this issue and enhance the reliability of LVLMs.
@@ 3.2 The Proposed CMI-VLD
(1) Calibrated Distribution Sampling for Text Modality. To optimize the text sequence $y$, Eq. (4) encourages us to construct an improved distribution $p_{c}$ to prioritize text tokens that maximize the difference between probabilities predicted with and without the visual input. However, directly applying this formulation yields unsatisfactory results since it can excessively penalize reasonable tokens in certain contexts. Inspired by [17, 9], we introduce a hyperparameter $\lambda$ to provide a more fine-grained control over the strength of the subtraction, which can be formally expressed as:
Figure 2: Overview of the proposed CMI-VLD decoding. At each timestep $t$, CMI-VLD mitigates hallucination by maximizing mutual dependency between the visual input and the ongoing response through the proposed vision-language purification. Specifically, the visual token purifier first incorporates current input tokens to predict an image mask $\mathcal{M}_{v}$, which filters out irrelevant visual tokens to enhance C-PMI. Based on the refined visual input, a text token distribution is correspondingly constructed to penalize hallucination-related text tokens and hence guide the next-token sampling to further strengthen the dependency on the visual input.
@@
(2) Visual Token Refinement for Visual Modality. Motivated by recent findings [37, 38] that many image tokens in LVLMs are redundant, we propose a visual token purification mechanism that enhances C-PMI by evicting tokens considered non-informative with respect to the given text. In this way, the LVLM can focus more on the most critical visual tokens for improved generation. Moreover, we also incorporate the model’s attention scores over the visual input to identify tokens that exert a stronger influence on the LVLM’s decisions. Given the query vectors $Q_{i}\in\mathbb{R}^{H\times n\times d_{k}}$ and key vectors $K_{i}\in\mathbb{R}^{H\times n\times d_{k}}$ at the $i$-th LVLM layer, where $H$ is the head number, $n$ is the current token number, and $d_{k}$ is the latent dimension, the total attention scores of an image $v$ is calculated as:
@@
where $\lambda$ is the aforementioned correction factor for distribution calibration and $\alpha$ is a non-negative hyperparameter balancing the influence of C-PMI and attention scores on visual token selection.
At decoding step $t$, we first optimize the upper subproblem by sampling $y_{t}$ from the distribution adjusted based on Eq. (5). To solve the lower subproblem, we then adaptively retain a proportion $\gamma$ of image tokens as the purified input to promote its relevance to the current textual context. Motivated by findings in [38] that token sparsification at the second layer of LVLMs yields optimal performance, we utilize attention scores from this layer ($i=2$) for visual purification and start the refinement accordingly. By alternately solving the two subproblems at each decoding step, the optimizer simultaneously samples text tokens that tightly align with current visual content and purifies informative visual tokens with strong relevance to the ongoing textual context, hence effectively amplifying C-PMI and reducing hallucination-related elements in the final response. For stable performance, we also incorporate the feature steering mechanism [39] into our implementation.
@@ 3.3 Visual Token Purifier for Visual Refinement
To address the outer subproblem in Eq. (7), an intuitive solution is to manually select image tokens that maximize the defined score for every decoding step. However, this requires repeatedly calculating token-wise scores and would incur substantial computational burdens compared to existing decoding-based approaches [17, 18]. To overcome this challenge, we propose a lightweight visual token purifier $\mathcal{P}(\cdot)$, which consists of only a few transformer blocks and MLP layers (see Appendix D for details) [37], to automatically filter visual tokens that benefit C-PMI maximization.
@@ 4.1 Experimental Setup
Models and Baselines. We align with [9] and choose four representative LVLMs for evaluation, including InstructBLIP [31], Shikra [3], LLaVA-1.5 on the 7B scale [4], and LLaVA-NeXT [32] on the 8B scale. We conduct a comprehensive evaluation of the proposed CMI-VLD on a range of state-of-the-art (SOTA) baselines, including Sampling (Top-p=1), Greedy, VTI [39], VCD [17], ICD [18], HALC [8], OPERA [19], SID [9], and VASparse [10]. Following SID, we implement the proposed method under both sampling and greedy decoding settings. It is worth noting that HALC, OPERA, and VASparse adopt the more flexible and stronger beam search strategy, which may raise fairness concerns as it retains a broader set of promising candidate paths during decoding. Nevertheless, our CMI-VLD still consistently outperforms these methods.
@@
Implementation Details. We set $\alpha$ = $1\times 10^{2}$ and $\beta$ = $5\times 10^{2}$ in Eq. (9) during the training of the purifier for all LVLMs. To initiate the refinement process, we adopt Layer $i=2$ for LLaVA-1.5, Shikra, and LLaVA-NeXT and $i=4$ for InstructBLIP. In Sec. 4.3, we explore the contrastive strength $\lambda$ ranging from 0 to $0.9$. For all experiments, we set max new tokens as 512 for evaluation. Note that CMI-VLD is compatible with the feature steering mechanism in VTI [39], which is then incorporated in our implementation for stable and enhanced performance. More details are in Appendix B.
4.2 Performance Evaluation
@@
POPE Evaluation. The POPE metric also focuses on object hallucinations by using a prompt "Is there a <object> in the image?" to query LVLMs for answering a yes/no question. We report the results of the accuracy and F1 score in Table 2. The quantitative results reveal that our method generally performs best across the three split datasets. Notably, in the POPE evaluation, where responses are typically short and follow fixed patterns such as "Yes, there is a <object> in the image.", the evaluation primarily hinges on the first one or few tokens (i.e., Yes or No)  [19]. Consequently, our CMI-VLD may not fully exhibit its effectiveness in this constrained setup, as it is designed to dynamically adjust decoding over the entire generation rather than concentrating solely on the initial tokens. Nevertheless, our method still achieves notable improvements over competitive baselines. Due to the page limit, results on more LVLMs are provided in Appendix C.
MME and MMBench Evaluations. Apart from the above benchmarks tailored for hallucination evaluation, we additionally test on two popular LVLM benchmarks, i.e., MME [7] and MMBench [45], to systematically analyze their various capability dimensions. MME provides a suite of fine-grained, image-grounded multiple-choice questions across various categories. We follow SID and report the overall perception score covering 10 sub-tasks, such as object existence and fine-grained recognition. MMBench is another large-scale bilingual benchmark consisting of over 3,000 curated multiple-choice questions. We compute the LVLM’s average score across 20 multimodal tasks, such as attributes, logical reasoning, and coarse/fine-grained perception, to comprehensively evaluate its capabilities. As observed, CMI-VLD not only reduces the hallucinated contents but also enhances diverse capabilities of MLLMs, bringing remarkable improvements over the default decoding methods. These results underscore CMI-VLD as a reliable and practical strategy for hallucination mitigation.
@@
Inference Time Analysis. Since our method introduces an additional visual token purifier for effective visual refinement, it is crucial to assess its influence on the overall computational efficiency. Following [20], we calculate the generation time per response based on LLaVA-1.5 to assess computing burdens. Figure 4 reveals that the proposed CMI-VLD achieves satisfactory decoding efficiency, introducing negligible computational overhead. This is primarily attributed to the lightweight architecture of the visual purifier and the removal of redundant visual tokens that would incur significant computational overhead, demonstrating that CMI-VLD effectively balances performance and efficiency.
4.3 Ablation Study
Next, we provide ablation analysis regarding several critical hyperparameters. More ablation analysis about the retaining ratio and the effectiveness of the proposed techniques is presented in Appendix C.
[…]